OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –(x + 4)?

OpenStudy (anonymous):

y - y1=m(x-x1)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

change it to y = mx + b start by distributing the negative

OpenStudy (anonymous):